Solve
The square root of x+1, Under the square root of 3=the square root of 3x-5

Answer by edjones(8007)

sqrt(3)/sqrt(x+1)=sqrt(3x-5)
3/(x+1)=3x-5
3=3x^2-2x-5
3x^2-2x-8=0
3x^2-6x+4x-8=0
3x(x-2)+4(x-2)=0
3x+4=0, x-2=0
x=2
.
Ed